OpenStudy (anonymous):

solving radical equations ^4 square root x-7=2 heeelp!

OpenStudy (anonymous):

\[\sqrt[4]{x-7=2}\]

OpenStudy (anonymous):

the answer is 3

OpenStudy (anonymous):

the square root of 9 which is what x-7=2 = so the square root of 9 is 3

OpenStudy (anonymous):

what is the process?

OpenStudy (anonymous):

cant be 3

OpenStudy (anonymous):

look carefully its not the square root but the 4th root

OpenStudy (anonymous):

is the 2 also under the root sign?

OpenStudy (anonymous):

I don't know how do this

OpenStudy (anonymous):

@midhun.madhu1987 do you know this?

OpenStudy (midhun.madhu1987):

Question unclear... = cannot be inside square root

OpenStudy (anonymous):

both side power of 4th

OpenStudy (anonymous):

x-7=2^4

OpenStudy (anonymous):

x=16-7=9